OpenStudy (anonymous):

evening, trying to solve for x: 2x-5=-13

OpenStudy (anonymous):

2x-5=-13 2x=-13+5 2x=-8 x=-8/2 x=-4

OpenStudy (anonymous):

plug -4 into the original equation to see if it does eqaul -13 2x-5=-13 2(-4)-5=-13 -8-5=-13 -13=13 so your answer is correct. x=-4

OpenStudy (anonymous):

thanks! that is what we found too, but was hungup on x being a negative #. thanks

OpenStudy (anonymous):

no problem!:DD